# EventBind: Learning a Unified Representation to Bind Them All for Event-based Open-world Understanding

### v4

@@ Abstract
In this paper, we propose EventBind, a novel and effective framework that unleashes the potential of vision-language models (VLMs) for event-based recognition to compensate for the lack of large-scale event-based datasets. In particular, due to the distinct modality gap with the image-text data and the lack of large-scale datasets, learning a common representation space for images, texts, and events is non-trivial. Intuitively, we need to address two key challenges: 1) how to generalize CLIP’s visual encoder to event data while fully leveraging events’ unique properties, e.g., sparsity and high temporal resolution; 2) how to effectively align the multi-modal embeddings, i.e., image, text, and events. Accordingly, we first introduce a novel event encoder that subtly models the temporal information from events and meanwhile generates event prompts for modality bridging. We then design a text encoder that generates content prompts and utilizes hybrid text prompts to enhance EventBind’s generalization ability across diverse datasets. With the proposed event encoder, text encoder, and image encoder, a novel Hierarchical Triple Contrastive Alignment (HTCA) module is introduced to jointly optimize the correlation and enable efficient knowledge transfer among the three modalities. We evaluate various settings, including fine-tuning and few-shot on three benchmarks and our EventBind achieves new state-of-art accuracy compared with the previous methods, such as on N-Caltech101(+5.34% and +1.70%) and N-Imagenet(+5.65% and +1.99%) with fine-tuning and 20-shot settings respectively. Moreover, our EventBind can be flexibly extended to the event retrieval task using text or image queries, showing plausible performance. Project homepage: https://vlislab22.github.io/ECLIP/.
@@
Multi-modal, Event camera, Object recognition, Retrieval
@@ 3 Proposed Approach
Figure 2: Overview of our proposed EventBind, which comprises of image encoder pre-aligned with text encoder and the proposed event encoder. It takes the image(optional), text, and event frames as input, generating the image embeddings $f^{i}$, text embeddings with the event content and image content $f^{e,i},f^{t,i}$, and event embeddings $f^{i}$. Then all output embeddings are aligned in the HTCA module.
@@
Consequently, we design our EventBind consisting of four major components, as depicted in Fig. 2: 1) the event encoder, responsible for extracting dense temporal and sparse spatial information from events and bridging the modality gap (Sec. 3.1); 2) the text encoder (Sec. 3.2) incorporates the event and image content and pre-trained generalizing knowledge from CLIP to generate the text embeddings for enhance performance; 3) the original CLIP’s image encoder for semantically aligned image embeddings  [^1] As the image encoder in E-CLIP remains identical to CLIP, we do not elaborate on it in the following subsection. Note that the image input is optional.; 4) The hierarchical triple contrastive alignment (HTCA) module (Sec. 3.3) jointly optimizes the correlation among the three distinct modalities, forming a unified embedding space among image, text, and event. In the following sections, we describe these technical components.
@@ 3.1 Event Encoder
Event Prompts Generation. Based on the event temporal modeling, the unique temporal correlation among event frames is captured. Since fine-tuning the pre-trained image encoder for processing event data may not guarantee good performance because of the modality gap and smaller event dataset size compared to the web-scale image dataset, we propose the event prompts to bridge the modality gaps among the image, text, and event modalities without introducing excessive parameters, which may cause catastrophic forgetting [24]. Specifically, we insert a small number of learnable parameters at the input of each encoder layer. The proposed event prompts $P_{e}$ are learnable vectors inserted at the output token sequence between the original class token and other message tokens (see orange rectangles in Fig. 3). After event temporal modeling and event prompts generation, we concatenate the original message token $Z^{s,t}$, event prompts $P^{s,t}_{e}$ and cross-frame prompts $x_{cf}^{s,t}$ and send it into a standard self-attention layer as:
@@ 3.3 Hierarchical Triple Contrastive Alignment (HTCA)
With the extracted event embeddings $f^{e}$(Sec. 3.1), image embeddings $f^{i}$, and text embeddings with event content prompt $f^{t,e}$(Sec. 3.2), it’s still an open question to bridge the modality gap among the three distinct modalities i.e., image, text, and event, and effectively align them. A simple objective is to align $f^{e}$ with $f^{t}$ from scratch. However, the relatively small event dataset and limited vocabulary amount may hinder reliable representation and plausible few-shot performance. To this end, we propose the Hierarchical Triple Contrastive Alignment (HTCA) module, aiming at aligning the event, image, and text modalities jointly and enhancing model performance based on semantic consistency.
Multi-modality Triple Alignment We first minimize the contrastive loss $L(f^{i},f^{e})$ between the event embeddings $f^{e}$ and image embeddings $f^{i}$, as well as aligning $L(f^{e},f^{t,e})$ the event embeddings $f^{t,e}$ with text embeddings with event content prompts $f^{t,e}$ by utilizing event-image and event-text pairs. We ignore the image-text pair because we empirically observe that altering the pre-trained image-text alignment reduces performance. Specifically, the contrastive loss is obtained by averaging the whole mini-batch among each pair of modalities $M_{1}$ and $M_{2}$, which can be formulated as follows:
@@ 4.1 Experimental Setup and Datasets
Dataset. We use three public image-event-text paired datasets for our experiments: N-ImageNet [25] with ImageNet-1K [12], N-Caltech101 [16] with Caltech101 [40], and N-MNIST [40] with MNIST [13]. N-ImageNet [25] is derived from the ImageNet-1K dataset, where the RGB images are displayed on a monitor and captured by a moving event camera. It includes 1,781,167 event streams with 480 × 640 resolution across 1,000 unique object classes. N-Caltech101 [16] contains event streams captured by an event camera in front of a mobile 180 × 240 ATIS system[41] with the LCD monitor presenting the original RGB images in Caltech101. There are 8,246 samples comprising 300 ms in length, covering 101 different types of items. N-MNIST is created by displaying a moving image from the MNIST dataset on the ATIS system with the LCD monitor. It contains 70,000 event data samples covering 10 handwritten numbers from 0 to 9. We split N-ImageNet and N-MNIST officially. N-Caltech101 is split 4:1 for training and validation. The same validation set is used for zero-shot, few-shot, and fine-tuning. Few-shot data are randomly selected and consistent. See more details on experimental settings in the Suppl.
@@
Inference and Task Settings. Event recognition indicates the classification of event data, which is a fundamental task in event-based vision. As shown in Fig. 1, text class names are input into the text encoder to generate text embeddings $f^{t,e}$. The event data is encoded to produce the event embeddings $f^{e}$, and then the softmax function is applied to the final probability map generated by the multiplication between $f^{t,e}$ and $f^{e}$. For open-world event recognition, the text classes are arbitrary rather than restricted to the training classes. In fine-tuning and few-shot event recognition scenarios, the former employs the entire training dataset, while the latter uses only a handful of training examples for each category during the training phase.
@@ 4.2 The Results for Event Recognition Task
Few-shot Setting. We compare our EventBind with the existing method, EventCLIP [55], on the N-Caltech101 and N-ImageNet datasets. To ensure comparison consistency, the ViT-L-14 backbone is utilized as EventCLIP stated. In Tab. 2, EventBind beats EventCLIP in few-shot cases, with gains of 0.92%, 1.31% and 1.7% for 5-shot, 10-shot, and 20-shot on N-Caltech101. Remarkably, our 20-shot accuracy outperforms the previous state-of-the-art (SOTA) [35], which was trained on the full dataset, while EventBind uses just 46.09% of the data from the N-Caltech101 dataset. For the N-ImageNet dataset, EventBind also outperforms EventCLIP with 2.78%, 2.73%, and 3.99% accuracy improvements for 5-shot, 10-shot, and 20-shot settings, showcasing its superiority in low-shot scenarios.
@@ 4.3 Ablation Study
Table 7: Ablation study on training configuration & event frame representation. Setting Top-1 Accuracy (%) N-Caltech101 N-MNIST Frozen CLIP 86.69 97.11 Fitune CLIP 94.08 99.27 Gray-scale 92.18 97.61 RGB 94.08 99.27 Table 8: Ablation study on different hand-crafted text prompts. Hand-crafted text prompts Top-1 Accuracy (%) N-Caltech101 N-MNIST ‘A photo of a’ 93.06 98.53 ‘A sketch image of a’ 93.89 99.20 ‘A point cloud image of a’ 92.71 98.49 ‘An event frame of a’ 92.41 98.87 ‘A drafted image of a’ 94.08 99.27
[…]

### v6

# Project Page: https://vlislab22.github.io/EventBind/

@@ Abstract
In this paper, we propose EventBind, a novel and effective framework that unleashes the potential of vision-language models (VLMs) for event-based recognition to compensate for the lack of large-scale event-based datasets. In particular, due to the distinct modality gap with the image-text data and the lack of large-scale datasets, learning a common representation space for images, texts, and events is non-trivial. Intuitively, we need to address two key challenges: 1) how to generalize CLIP’s visual encoder to event data while fully leveraging events’ unique properties, e.g., sparsity and high temporal resolution; 2) how to effectively align the multi-modal embeddings, i.e., image, text, and events. Accordingly, we first introduce a novel event encoder that subtly models the temporal information from events and meanwhile generates event prompts for modality bridging. We then design a text encoder that generates content prompts and utilizes hybrid text prompts to enhance EventBind’s generalization ability across diverse datasets. With the proposed event encoder, text encoder, and image encoder, a novel Hierarchical Triple Contrastive Alignment HTCA module is introduced to jointly optimize the correlation and enable efficient knowledge transfer among the three modalities. We evaluate various settings, including fine-tuning and few-shot on three benchmarks and our EventBind achieves new state-of-art accuracy compared with the previous methods, such as on N-Caltech101 (+5.34% and +1.70%) and N-Imagenet (+5.65% and +1.99%) with fine-tuning and 20-shot settings respectively. Moreover, our EventBind can be flexibly extended to the event retrieval task using text or image queries, showing plausible performance.
@@
Event Camera Multi-modal Object Recognition
@@ 3 Proposed Approach
Figure 2: Overview framework of EventBind, which consists of image encoder pre-aligned with text encoder and the proposed event encoder. It takes image(optional), text, and event as input, generating the image embeddings $f^{i}$, event embeddings $f^{e}$ and text embeddings $f^{t,e},f^{t,i}$. Then all output embeddings are aligned in the HTCA module to establish a unified representation space.
@@
Consequently, we design our EventBind consisting of four major components, as depicted in Fig. 2: 1) the event encoder, responsible for extracting dense temporal and sparse spatial information from events and bridging the modality gap (Sec. 3.1); 2) the text encoder (Sec. 3.2) incorporates the event and image content and pre-trained generalizing knowledge from CLIP to generate the text embeddings for enhance performance; 3) the original CLIP’s image encoder for semantically aligned image embeddings  [^1] As the image encoder in EventBind remains identical to CLIP, we do not elaborate on it in the following subsection. Note that the image input is optional.; 4) The hierarchical triple contrastive alignment (HTCA) module (Sec. 3.3) jointly optimizes the correlation among the three distinct modalities, forming a unified embedding space among image, text, and event. In the following sections, we describe these technical components.
@@ 3.1 Event Encoder
Event Prompts Generation. Based on the event temporal modeling, the unique temporal correlation among event frames is captured. Since fine-tuning the pre-trained image encoder for processing event data may not guarantee good performance because of the modality gap and smaller event dataset size compared to the web-scale image dataset, we propose the event prompts to bridge the modality gaps among the image, text, and event modalities without introducing excessive parameters, which may cause catastrophic forgetting [28]. Specifically, we insert a small number of learnable parameters at the input of each encoder layer. The proposed event prompts $P_{e}$ are learnable vectors inserted at the output token sequence between the original class token and other message tokens (see orange rectangles in Fig. 3). After event temporal modeling and event prompts generation, we concatenate the original message token $Z^{s,t}$, event prompts $P^{s,t}_{e}$ and cross-frame prompts $x_{cf}^{s,t}$ then send it into a standard self-attention layer as:
@@ 3.3 Hierarchical Triple Contrastive Alignment (HTCA)
With the extracted event embeddings $f^{e}$ (Sec. 3.1), image embeddings $f^{i}$, and text embeddings with event content prompt $f^{t,e}$ (Sec. 3.2), it’s still an open question to bridge the modality gap among the three distinct modalities i.e., image, text, and event, and effectively align them. A simple objective is to align $f^{e}$ with $f^{t}$ from scratch. However, the relatively small event dataset and limited vocabulary may hinder reliable representation and plausible few-shot performance. To this end, we propose the Hierarchical Triple Contrastive Alignment (HTCA) module, aiming at aligning the event, image, and text modalities jointly and enhancing model performance based on semantic consistency.
Multi-modality Triple Alignment We first minimize the contrastive loss $L(f^{i},f^{e})$ between the event embeddings $f^{e}$ and image embeddings $f^{i}$, as well as minimizing $L(f^{e},f^{t,e})$ between the event embeddings $f^{t,e}$ and text embeddings with event content prompts $f^{t,e}$. We ignore the image-text pair because we empirically observe that altering the pre-trained image-text alignment reduces performance. Specifically, the contrastive loss is obtained by averaging the whole mini-batch among each pair of modalities $M_{1}$ and $M_{2}$, which can be formulated as follows:
@@ 4.1 Experimental Setup and Datasets
Dataset. We use three public image-event-text paired datasets for our experiments: N-ImageNet [29] with ImageNet-1K [13], N-Caltech101 [19] with Caltech101 [46], and N-MNIST [46] with MNIST [14]. We split N-ImageNet and N-MNIST officially. N-Caltech101 is split 4:1 for training and validation. The same validation set is used for zero-shot, few-shot, and fine-tuning. Few-shot data are randomly selected and consistent. See more details on experimental settings in the Suppl.
@@
Inference and Task Settings. Event recognition indicates the classification of event data, which is a fundamental task in event-based vision. As shown in Fig. 1, text class names are input into the text encoder to generate text embeddings $f^{t,e}$. The event data is encoded to produce the event embeddings $f^{e}$, and then the softmax function is applied to the final probability map generated by the multiplication between $f^{t,e}$ and $f^{e}$. In fine-tuning and few-shot event recognition scenarios, the former employs the entire training dataset, while the latter uses only a handful of training examples for each category during the training phase.
@@ 4.2 The Results for Event Recognition Task
Few-shot Setting. We compare EventBind with existing methods, EventCLIP [64]. To ensure comparison consistency, the ViT-L-14 backbone is utilized as EventCLIP stated. In Tab. 2, EventBind beats EventCLIP in few-shot cases, with gains of 0.92%, 1.31% and 1.7% for 5-shot, 10-shot, and 20-shot on N-Caltech101. Remarkably, our 20-shot accuracy outperforms the previous state-of-the-art (SOTA) [40], which was trained on the full dataset, while EventBind uses just 46.09% of the data from the N-Caltech101 dataset. For the N-ImageNet dataset, EventBind also outperforms EventCLIP with 2.78%, 2.73%, and 3.99% accuracy improvements for 5-shot, 10-shot, and 20-shot settings, showcasing its superiority in low-shot scenarios.
@@ 4.3 Ablation Study
Table 3: Ablation study on EventBind’s key components.
Ablation Settings Top-1 Accuracy (%)
HTCA Text prompt Event encoder N-Caltech101 N-MNIST
@@
[…]
Table 4: Ablation study on the HTCA module. ‘_‘ highlights the EventBind performance in the absence of image inputs.
Ablation Settings Top-1 Accuracy(%)
$L(f^{e},f^{i})$ $L(f^{e},f^{t,e})$ $L(f^{t,i},f^{t,e})$ N-Caltech101 N-MNIST
[…]
Table 5: Ablation study on event encoder. $P_{e}$ means event prompts.
Setting Top-1 Accuracy (%)
$P_{e}$ Event Temporal modeling N-Caltech101 N-MNIST
[…]
Table 6: Model parameter analysis.
Backbone ViT-L-14 ViT-B-32 ViT-B-16
EventCLIP[64] 890 MB - 335 MB
EventBind (w.o Image Encoder) 1372 MB 474 MB 471 MB
EventBind (w.t. Image Encoder) 1952 MB 641 MB 635 MB
@@
Table 7: Ablation study on training configuration & event frame representation.
Setting Top-1 Accuracy (%)
N-Caltech101 N-MNIST
Frozen CLIP 86.69 97.11
Fitune CLIP 94.08 99.27
Gray-scale 92.18 97.61
RGB 94.08 99.27
Table 8: Ablation study on different hand-crafted text prompts.
Hand-crafted text prompts Top-1 Accuracy (%)
N-Caltech101 N-MNIST
‘A photo of a’ 93.06 98.53
‘A sketch image of a’ 93.89 99.20
‘A point cloud image of a’ 92.71 98.49
‘An event frame of a’ 92.41 98.87
‘A drafted image of a’ 94.08 99.27
Table 9: Abaltion study on the content prompts generation.
Abaltion settings Top-1 Accuracy (%)
$\sigma_{i}$ $\sigma_{e}$ N-Caltech101 N-MNIST
[…]
Table 10: Ablation study on length of event modality prompts.
$n_{e}$ Top1 Accuracy
N-Caltech101 N-MNIST
4 93.27 97.95
8 93.89 98.73
16 94.08 99.27
Table 11: Ablation study on length of text learnable prompt.
$n_{l}$ Top1 Accuracy
N-Caltech101 N-MNIST
4 93.33 98.10
8 93.75 98.83
16 94.08 99.27
@@ 4.4 Extension Experiment: Event Retrieval
Figure 5: Visual examples for the event retrieval task.
EventBind can be expanded to image-to-event and text-to-event retrieval tasks. See more details about event retrieval and results discussions in Suppl. We initially perform the event retrieval task by employing both text and image queries. In Fig. 5, we present the Top-5 event data retrieved from the N-Caltech101 dataset using randomly selected images or class texts as queries. All the retrieved event data exhibit a very high degree of similarity to the input text query, showcasing the effectiveness of our proposed EventBind framework.
5 Conclusion and Future Work
Acknowledgements
This paper is supported by the National Natural Science Foundation of China (NSF) under Grant No. NSFC22FYT45, the Guangzhou City, University and Enterprise Joint Fund under Grant No.SL2022A03J01278, and Guangzhou Fundamental and Applied Basic Research (Grant Number: 2024A04J4072)
Appendix
A. Preliminary of CLIP
The Contrastive Language-Image Pre-training (CLIP) [50] model is a prominent vision-language model that combines an image encoder (ViT[17] or ResNet[25]), with a text encoder based on the Transformer [57] architecture. CLIP focuses on generating visual and text embeddings and employs a contrastive loss to align these embeddings in a unified feature space. Notably, CLIP’s exceptional transferability is attributed to its pretraining on a large-scale dataset of more than four million image-text pairs.
For CLIP-based recognition, it usually employs a simple yet efficient method that takes N hand-crafted prompts like ‘a photo of a $[cls]$’ as the input for the text encoder to obtain the D-dimensional textual embedding $f^{t}\in\mathbb{R}^{N\times D}$, where $[cls]$ is the class name and N is the number of classes of the downstream dataset. Meanwhile, for a given image, the visual embedding $f^{i}\in\mathbb{R}^{1\times D}$ is obtained by the visual encoder. Then, the recognition logits $p\in\mathbb{R}^{N}$ can be obtained by calculating the similarity of text and visual embeddings by multiplication:
$p=\mathrm{Softmax}(f^{i}(f^{t})^{T}),$ (6)
where $\mathrm{Softmax}(.)$ denotes the Softmax function and p represents the predicted probability of the N classes. Finally, the highest scores of the logits p is regarded as the final prediction P:
$P=\arg\max_{i\in\mathcal{N}}p_{i}$ (7)
B. Implementation Details
Event Frame-like Representation Event cameras capture object movement by detecting temporal independence and recording pixel-level brightness changes. The event stream, denoted as $\mathcal{E}={e_{i}(x_{i},y_{i},t_{i},p_{i})}$, reflects the brightness change $e_{i}$ of a pixel at the timestamp $t_{i}$, with coordinates $(x_{i},y_{i})$, and polarity $p_{i}\in{1,-1}$. Here, 1 and -1 represent the positive and negative intensity change of brightness, respectively. We convert the event stream into a sequence of frames. First, we normalize the length of the event stream $\mathcal{E}$ to a fixed amount $P$ using zero-padding or taking the first $P$ events. Then, we group every $Q$ consecutive event data in the normalized event stream $\mathcal{E}^{{}^{\prime}}$ to obtain event parts ′′$\mathcal{E}^{{}^{\prime\prime}}\in\mathbb{R}^{T\times 4\times P}$, where $T=P/Q$ and denotes the number of event frames.
We then transform these event parts into histograms $h\in\mathbb{R}^{T\times H\times W\times 2}$ by counting the amount of positive and negative events per pixel, where $H$ and $W$ represent the height and width of an event frame, respectively. Finally, to attain the grid-like 3-channel input $I_{e}\in\mathbb{R}^{T\times H\times W\times 3}$, we colorize the $t$th event histogram $h_{t}$ by multiplying the predefined red color map $[0,255,255]$ and blue color map $[255,255,0]$ with the positive and negative event histogram respectively and merge them by pixel-wise addition, which can be formulated as follow:
$I_{e,t}=h_{t}^{p}[0,255,255]^{T}+h_{t}^{n}[255,255,0]^{T},$ (8)
where the $I_{e,t}$ represents the event input for the $t$th event frame and $t=0,1,...,T$; $h_{t}^{p}$ and $h_{t}^{n}$ denotes the positive and negative event histogram respectively. Finally, the generated event frame input is resized into the 224 $\times$ 224 resolution for adapting to the ViT setting, generating the final event input $I_{e}^{{}^{\prime}}\in\mathbb{R}^{T\times 224\times 224\times 3}$.
The above grid-like event representation is simple yet effective since the generated event image resembles the edge image, whose data distribution is much closer to the pre-trained natural images so that the transfer pressure is erased and the modality gap is bridged.
Backbone N-Imagnet N-MNIST N-Caltech101
ViT-B-16 300,000 300 150,000
ViT-B-32 300,000 400 150,000
ViT-L-14 400,000 1,000 200,000
Table 12: The settings of aggregated event counts per frame $N$ for different backbones on three datasets.
Additional Dataset Settings N-ImageNet [29] is derived from ImageNet-1K dataset, where the RGB images are displayed on a monitor and captured by a moving event camera. It includes 1,781,167 event streams with 480 × 640 resolution across 1,000 unique object classes. N-Caltech101 [19] contains event streams captured by an event camera in front of a mobile 180 × 240 ATIS system[49] with the LCD monitor presenting the original RGB images in Caltech101. There are 8,246 samples comprising 300 ms in length, covering 101 different types of items. N-MNIST is created by displaying a moving image from the MNIST dataset on the ATIS system with the LCD monitor. It contains 70,000 event data samples covering 10 handwritten numbers from 0 to 9.
Additional Experimental Settings We use "A drafted image of a [CLS]" as the hand-crafted text prompts template. The Pytorch [48] framework serves as the foundation for all experiments. The initial learning rates are set to 1e-5 for the N-Caltech101 dataset and 1e-6 for the N-Imagenet and N-MNIST datasets. The weight decay is 2e-4. CosineAnnealingLR [42] learning rate schedule is utilized, and the minimal learning rate is 1e-8. All few-shot and fine-tuning experiments are trained for 30 epochs with Adam [30] optimizer. Unless specified otherwise, the ablation study is conducted on the N-Caltech101/Caltech101 datasets utilizing ViT-B-16 [16] image encoder and the Transformer-based text encoder [57] as the backbone. The event encoder is initialized with the CLIP image encoder’s pre-trained weights and is fine-tuned during training. We choose the aggregated event counts per frame $N$ based on the best EventBind’s zero-shot performance. Tab. 12 represents the value of $N$ set for different backbones on three datasets.
Ablation settings Base New
Learnable Hand-crafted N-Caltech101 N-MNIST N-Caltech101 N-MNIST
[…]
Table 13: Ablation study on hybrid text prompts module.
C. Additional Experiment Result
The Effectiveness of Hybrid Text Prompts We ablated the hybrid text prompts to evaluate their impact on fine-tuning performance and generalization ability. The experiment follows the setting proposed in [75], where the entire dataset is divided into the base and new datasets, each containing half of the object classes. For each ablation setting, the model is fine-tuned on the base set without being exposed to half of the event classes in the new set. The evaluation is then performed on the new set to assess the generalization ability.
As shown in Tab. 13, the hand-crafted text prompts achieve an impressive 72.55% Top-1 accuracy on the new set for the N-Caltech101 dataset, showcasing its remarkable zero-shot ability. In contrast, while learnable text prompts boost fine-tuning performance on the base set, they lead to decreased accuracy on the new set due to limited generalization capabilities. In contrast, our proposed hybrid text prompts module, which integrates both learnable and manually-crafted text prompts, attains the highest fine-tuning accuracy of 91.23% while exhibiting a smaller decrease in zero-shot performance relative to solely utilizing single hand-crafted text prompts. (-10.70% V.S -4.53% for N-Caltech101). For the N-MNIST dataset, it’s expected that the model exhibits low accuracy on the base dataset with hand-crafted text prompts. This stems from the subpar image recognition capability of the original CLIP on N-MNIST, where zero-shot CLIP results are 10% inferior to those of the linear probe on ResNet50 for MNIST. [50]). Our model, equipped with the hybrid text prompts module, secures top-one accuracy on both the base dataset (93.56%) and the new dataset (98.86%) for N-MNIST, underscoring the effectiveness of our proposed module.
EventBind (scratch) EventCLIP (scratch)
Setting N-Calthech101 N-MINIST N-Calthech101 N-MINIST
0-shot 61.80 56.81 58.82 48.72
1-shot 74.96 74.64 70.53 74.62
2-shot 79.78 82.89 77.71 82.78
Table 14: The zero-shot and few-shot results.
Model E2VID[52] SSL-E2VID[47] Wang et al.[59] Ev-LaFOR[12] EventBind(Ours)
Top1 Accuracy 61.7 25.3 48.5 85.56 86.02
Table 15: Open-vocabulary results on N-Caltech101.
Zero-shot, 1-shot, 2-shot Results Compared with EventCLIP. Since the dataset split of EventCLIP for 1-shot and 2-shot is unavailable, we evaluated it on our dataset split with ViT-L-14 backbones. As shown in Tab. 14, our EventBind achieves better zero-shot, 1-shot, and 2-shot performance than EventCLIP.
Open-vocabulary Recognition Result For open-vocabulary event recognition, the text classes are arbitrary rather than restricted to the training classes. We follow the dataset split based on its open-source repository. The results are in Tab. 15. Our EventBind achieves the SoTA performance of 86.02%, proving its superiority for open-vocabulary recognition.
(a) (a) N-Caltech101
(b) (b) N-Imagenet
(c) (c) N-MNIST
Figure 6: The Accuracy curves with three ViT backbones on the N-MINIST, N-Caltech101, and N-Imagenet datasets.
Accuracy Curve We present the training accuracy curves trained by employing three ViT backbones on the N-MINIST, N-Caltech101, and N-Imagenet datasets in Fig. 6. As the number of training epochs increases, the accuracy of the model increases steadily, demonstrating the stability and reproducibility of EventBind.
Text2Event Imgae2Event
Setting Recall@1 Recall@5 Recall@10 Recall@1 Recall@5 Recall@10
0-shot 78.22 88.12 91.09 79.40 90.34 93.58
1-shot 81.19 95.05 98.02 82.75 96.04 99.01
2-shot 90.10 99.01 100.00 88.31 98.02 99.01
5-shot 97.03 100.00 100.00 90.97 99.01 100.00
10-shot 98.02 100.00 100.00 94.73 99.01 100.00
20-shot 99.01 100.00 100.00 96.70 100.00 100.00
Fine-tune 100.00 100.00 100.00 96.88 99.60 99.95
Table 16: Event Retrieval Results on N-Caltech101 Dataset with Zero-shot, Few-shot, and Fine-tuning Settings.